If the sum of the three consecutive number is 12 and x is represents the smallest number, what is the value of x?
pls answer this with complete solution.

Answer by algebrahouse.com(1659)

"If the sum of the three consecutive number is 12 and x is represents the smallest number, what is the value of x?"

x = 1st consecutive number
x + 1 = 2nd consecutive number
x + 2 = 3rd consecutive number

x + x + 1 + x + 2 = 12 {the sum of the three numbers is 12}
3x + 3 = 12 {combined like terms}
3x = 9 {subtracted 3 from both sides}
x = 3 {divided both sides by 3}
www.algebrahouse.com